the sum of numenator and denominator of a fraction is 17. if three is added to the numinator, the value of fraction will be 1. what is the fraction?
Answer by richwmiller(17219)

a+b=17
(a+3)/b=1
a=7 b=10
PS What is a numenator or even numinator? Was ist ein numenator?